OpenStudy (anonymous):

Algebra 1 9th grade HELP PLEASE!!!! 2x + y = 11 3x + 4y = 24

OpenStudy (anonymous):

By setting the equations equal to each other, we can eliminate a variable. Multiplying the top by -4 gives us: -8x -4y = -44 3x+4y = 24 now add the two: -5x=-20 x=4 plug back in: 2(4)+y=11 8+3 =11 3(4)+4(3) =12 + 12 =24! Check.

OpenStudy (anonymous):

Forgot to add: x=4, y=3.

OpenStudy (anonymous):

Can you do it with substitution pwz :3??

OpenStudy (anonymous):

equation 1: 2x + y = 11 equation 2: 3x + 4y = 24 equation 1: y = 11-2x plug it into equation 2 3x+4(11-2x) = 24 3x +44 -8x = 24 -5x = -20 x = 4 plug back into equation 1, now you know what x is. y = 11-2x post what you get for y.

OpenStudy (anonymous):

@tomo i got 3 :D

OpenStudy (anonymous):

yes that is correct.

OpenStudy (anonymous):

@tomo Thank youuuuu <3!!!

OpenStudy (anonymous):

your welcome